Verify the following identity by using an angle sum identity:
cos(2x) = 1 – 2(sin^2x). Hint (2x = x + x)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Verify the following identity by using an angle sum identity:
cos(2x) = 1 – 2(sin^2x). Hint (2x = x + x)
----------
Wow, helpful hint.
-----
cos(A + B) = cos(A)cos(B) - sin(A)sin(B)

Answer by stanbon(75887)

Verify the following identity by using an angle sum identity:
cos(2x) = 1 – 2(sin^2x). Hint (2x = x + x)
--------------------
cos(2x) = cos(x+x) = cos^2(x) -sin^2(x) = 1-sin^2(x)-sin^2(x) = 1-2sin^2(x)
=====================
Cheers,
Stan H.